A ladder is resting against a wall. The top of the ladder touches the wall at a height of 15 feet. Find the length of the ladder if the length is 5 feet more than its distance from the wall.
Answer by rfer(16322)

c^2=a^2+b^2
c^2=5^2+15^2
c^2=25+225
c^2=250
c=15.8 ft long